OpenStudy (erinweeks):

Solve the inequality. Graph the solution. 2|x + 1/3| < 9 HELP SOLVE THIS

OpenStudy (anonymous):

\[\frac{ 2\left| x+1/3 \right| }{ 2 }<\frac{ 9 }{ 2 } \rightarrow \left| x+1/3 \right|<\frac{ 9 }{ ?}\] Your solutions are: \[x+1/3<9/2\] and \[x+1/3<-9/2\]

OpenStudy (anonymous):

? should be a 2.

OpenStudy (erinweeks):

imma send the graphs because thats not corect

OpenStudy (erinweeks):

nevermind i cant but A. -4 5/6 < x < 4 1/6 B. -4 1/3 < x < 4 1/3 C. x< -4 1/3 or x> 4 1/3 D. x < -4 5/6 or x > 4 1/6

OpenStudy (anonymous):

Ugh. Fractions are nasty. A.

OpenStudy (erinweeks):

thank you. and i know i hate doing these kind of problems.

OpenStudy (anonymous):

I just very much dislike fractions...even though they are now part of my coursework. BOOOO.